You tune an organ such that a particular note has a fundamental frequency of 314Hz . If you play that note along with another note which is an equal temperament fifth (7 half steps) above it, what will be the beat rate between the 3rd harmonic of the lower note and the 2nd harmonic of the upper note?
Answer #1

fundamental frequency or lower note = 314 Hz

frequency of higher note = 314 + 5(314/12) = 445 Hz

3rd harmonic of lower note = 314 3 = 942 Hz

2nd harmonic of higher note = 445 2 = 890

beat rate = 942 - 890 = 52 Hz